Prove that
a) Any Open ball is an open set
b) Any closed ball is closed set.
1
Expert's answer
2020-07-02T04:47:00-0400

1. ANY OPEN BALL IS AN OPEN SET

proof: Let B(x0,r) be an open ball with center x0 and radius r in a metric space (X,d)

Let y€B(x0,r)

Define r1 =r-d(x0,y)

We claim that B(y,r1) belongs to B(x0,r)

To see this let z€B(y,r1)

Then

d(z,x0) < d(z,y) +d(y,x0) < r1+(r-r1)

So z€B(x0,r)

So B(y,r1) belongs to B(x0,r)

Hence B(x0,r) is an open set.


2. ANY CLOSE BALL IS A CLOSED SET

proof: Let B̅(x,r) be a closed ball. We prove that B̅'(x,r)= C(say) is an open ball.

Let y€C then d(x,y)>r

Let r1=d(x,y) then r1>r and take r2=r1-r

Consider the open ball B(y,r2/2) we prove that B(y,r2/2) belongs to C.

For this let z€B(y,r2/2) then d(z,y)<r2/2

By triangular inequality

d(x,y)<d(x,z)+d(z,y)

d(x,y)<d(z,x)+d(z,y)

d(z,x)>d(x,y)-d(z,y)

d(z,x)>r1-r2/2 = (2r1-r2)/2

d(z,x)> (2r1-r1+r)/2 , r2=r1-r

d(z,x)>(r1+r)/2

d(z,x)>(r+r)/2 , r1>r

d(z,x)>r


z does not belong to B̅(x,r)

This shows that z€C

B(y,r2/2) belongs to C.

Hence C is an open set. So B̅(x,r) is closed.